# Reliable Grid Forecasting: State Space Models for Safety-Critical Energy Systems

### v5

@@ 4.3 Bias-Constrained Probabilistic Objective
Practically, we use these constrained objectives during fine-tuning to learn a calibrated distribution while preventing trivial “safety” via inflation. We then use (via ) to select an operationally conservative schedule or reserve policy from the learned distribution, and we report Bias/OPR jointly with tail-risk metrics to ensure improvements are not attributable to trivial inflation. Appendix F describes how we empirically estimate from CAISO market data [20].
5 Experimental Setup
5.1 Dataset
@@
Preprocessing. Load values are normalized using z-score standardization computed only on training data to prevent information leakage. Temporal features include hour-of-day (0–23) and day-of-week (0–6), encoded as sinusoidal embeddings following standard practice [6]. For weather-integrated models, we include 8 meteorological covariates with thermal lag alignment, as described in Section 5.3.
5.2 CAISO Market Context and Benchmarks
@@
5.3 Thermal Lag Parameter Determination
@@
Evaluation Protocol (Walk-Forward Backtest). We evaluate using rolling-origin walk-forward backtesting over Nov 2023–Nov 2025. Starting with an initial 180-day training history, we refit every 90 days using an expanding training window; each refit uses a 30-day validation window immediately preceding the cutoff for early stopping. We then evaluate one forecast per day (stride 24h) over the subsequent block until the next cutoff (48h horizon). All reported metrics in Tables 1–2 are computed on the concatenated set of walk-forward timestamps (477 evaluation windows for CA ISO-TAC).
5.4 Models Under Comparison
We evaluate three Mamba-based architectures against established baselines, enabling direct comparison across parameter efficiency and accuracy.
Proposed Mamba Architectures.
@@ 5.4 Models Under Comparison
Baselines.
LSTM [22]: 2-layer LSTM (0.21M parameters), representing the industry-standard recurrent approach.
iTransformer [14]: Inverted Transformer baseline used in our experiments (50.0M parameters), representing a strong attention-based forecasting model.
Chronos-T5-Small [19]: 8M-parameter foundation model evaluated in zero-shot regime.
This comparison directly tests whether Mamba’s efficiency can match the accuracy of larger Transformer models. PowerMamba (2.5M parameters) is particularly compact, while S-Mamba and Mamba-ProbTSF (16.4M parameters) trade parameter count for expressive projection heads.
5.5 Model Hyperparameters
Following the experimental protocol of PowerMamba [9], all Mamba variants share consistent architectural hyperparameters: (embedding dimension), (state dimension), (convolution kernel), expansion factor 2, and 2 bidirectional encoder layers. Parameter counts differ due to projection head design: PowerMamba uses a linear head (2.5M total), while S-Mamba and Mamba-ProbTSF use MLP heads (16.4M total). Dropout rate is 0.1 across all models. Weather integration adds negligible parameters (1K for the weather embedding layer).
5.6 Training Protocol
@@
Fixed-split fine-tuning (loss-function ablation). For the loss-function ablation and bias/OPR constraints (Table 2, Table 3), we fine-tune weather-initialized checkpoints on a fixed split using the trainer framework: AdamW (lr ) with OneCycleLR, up to 150 epochs with patience 60, and multi-quantile heads (, weights ). Bias/OPR constraints are applied at the 24h lead time (; index 23 if using 0-based arrays for a 48h horizon forecast).
Reproducibility. All experiments run on NVIDIA RTX 5090 GPUs. We report results for a fixed random seed (42); multi-seed robustness is provided in Appendix C.
5.7 Evaluation Metrics
@@ 5.8 Statistical Analysis
6 Results
6.1 Multi-Horizon Performance (walk-forward)
Evaluation regime. This subsection reports results from the rolling-origin walk-forward backtest described in Section 5.3.
Table 1 presents multi-horizon MAPE. PowerMamba is the strongest Mamba variant at the 1-hour horizon, while iTransformer achieves the lowest MAPE at longer horizons (6h–24h) in our walk-forward evaluation. These accuracy rankings, however, do not fully determine operational tail risk (Reserve), motivating the grid-specific metrics reported later.
| Model | Params | 1h | 6h | 12h | 24h |
| S-Mamba | 16.4M | 5.49 0.12 | 3.86 0.02 | 3.40 0.47 | 8.03 0.18 |
| Mamba-ProbTSF | 16.4M | 5.23 0.04 | 3.91 0.19 | 3.30 0.08 | 7.89 0.17 |
| PowerMamba | 2.5M | 4.44 0.20 | 4.10 0.09 | 3.62 0.30 | 7.31 0.06 |
| LSTM | 0.21M | 7.85 0.20 | 5.24 0.13 | 9.46 0.32 | 9.99 0.13 |
| iTransformer | 50.0M | 4.64 0.10 | 3.36 0.05 | 3.10 0.21 | 6.85 0.08 |
| Chronos (zero-shot)† | 8.0M | 2.10 | 3.21 | 3.72 | 7.59 |
| †Zero-shot evaluation; no training or uncertainty estimation. | |||||
6.2 Error Analysis: The Impact of Weather on Forecast Reliability (walk-forward)
Evaluation regime. This subsection uses the walk-forward backtest forecasts (all lead times within the 48h horizon) described in Section 5.3.
@@ 6.2 Error Analysis: The Impact of Weather on Forecast Reliability (walk-forward)
6.3 Weather Integration Results (walk-forward)
Evaluation regime. This subsection reports walk-forward results on matched evaluation windows (Section 5.3).
Based on the error analysis, we integrated thermal-lagged weather covariates into the architectures. As shown in Figure 3, weather integration changes the forecast error distribution; improvements are not uniform across architectures and regions and should be reported on matched evaluation windows. Comprehensive 24-hour MAPE results across all TAC areas are provided in Appendix Table 7.
@@ 6.3 Weather Integration Results (walk-forward)
6.4 Probabilistic Calibration and Bias/OPR-Constrained Objectives (fixed split)
Evaluation regime. This subsection reports fixed-split fine-tuning results for loss-function ablations (Section 5.6); it is not directly comparable to the walk-forward results above.
@@ 6.4 Probabilistic Calibration and Bias/OPR-Constrained Objectives (fixed split)
Table 2 presents an operator-legible breakdown (CA ISO-TAC, 24h), explicitly reporting Bias24h and OPR alongside tail-risk Reserve. Three high-signal observations:
@@
| Model | Variant | Bias24h (MW) | Reserve | UPR (%) | OPR (%) |
| S-Mamba | MSE | +243.6 | 16.28 | 43.5 | 56.5 |
| Multi-Q | +405.4 | 13.65 | 34.0 | 66.0 | |
| +BiasCtrl | +507.6 | 12.75 | 31.2 | 68.8 | |
| PowerMamba | MSE | +370.9 | 12.91 | 38.4 | 61.6 |
| Multi-Q | +480.1 | 13.65 | 34.6 | 65.4 | |
| +BiasCtrl | +2027.1 | 7.42 | 6.2 | 93.8 | |
| Mamba- ProbTSF | MSE | +176.5 | 14.86 | 44.9 | 55.1 |
| Multi-Q | +355.2 | 14.17 | 36.2 | 63.8 | |
| +BiasCtrl | -325.9 | 17.45 | 64.3 | 35.7 | |
| iTransformer | MSE | +109.3 | 28.70 | 50.6 | 49.4 |
| Multi-Q | +1862.3 | 13.83 | 21.2 | 78.8 | |
| +BiasCtrl | +456.4 | 15.18 | 38.4 | 61.6 |
Effect of probabilistic calibration on tail events. Figure 5 compares (i) Weather MSE, (ii) WeatherMulti-Q, and (iii) WeatherMulti-Q+BiasCtrl on a single consistent evaluation split (same targets across variants). Multi-quantile calibration reduces large-error counts relative to Weather MSE, while adding Bias/OPR constraints partially trades off tail improvements for reduced inflation risk (Table 2).
6.5 NEMs (BTM Registry) Integration Results
Evaluation regime. This subsection reports walk-forward results on utilities where NEMs/registry features are available. We next evaluate whether static NEMs/registry-derived BTM features (e.g., installed PV and storage capacity) provide incremental predictive value on top of the weather-aligned baseline. This analysis is reported only for utilities where NEMs/registry features were available in our pipeline (PGE, SCE, SDGE, TIDC).
How NEMs features are integrated. NEMs/registry features are static or slow-moving, so we integrate them as an additional BTM input stream that is fused into the weather-integrated forecaster. To avoid physically irrelevant feature injection (e.g., PV capacity “explaining” nighttime net load), we apply a simple daylight prior (implemented as a time-of-day gate). When weather inputs are present, we further apply a lightweight context-conditioned modulation so the BTM contribution can vary smoothly by regime (e.g., daylight/irradiance conditions) rather than acting as an unconditional bias.
then fuse additively before the sequence encoder:
Figure 6 summarizes this “weather + NEMs” fusion path.
@@ 6.5 NEMs (BTM Registry) Integration Results
7 Discussion
7.1 Performance Comparison with CAISO Benchmarks
Table 4 provides contextual comparison with published CAISO operational forecasts and commercial alternatives.
@@ 7.1 Performance Comparison with CAISO Benchmarks
| LSTM Baseline | 6.49% | +42.7% | 0.21M | No |
| iTransformer | 5.37% | +18.0% | 50.0M | No |
@@
| iTransformer + Weather | 4.15% | -8.8% | 50.0M | Yes |
@@
The performance gains are particularly significant given the model efficiency: PowerMamba achieves competitive accuracy relative to published CAISO benchmarks with only 2.5M parameters—69% fewer than Chronos and 95% fewer than the iTransformer baseline used in our experiments. This efficiency enables real-time deployment on edge devices without sacrificing forecast quality.
7.2 Temperature, Tail Risk, and Operator Implications
Our error analysis shows a statistically significant but modest association between temperature and forecast error magnitude (), supporting the operational intuition that weather is a relevant covariate and that loss-function modifications cannot substitute for missing exogenous information. More importantly for operations, the walk-forward results reinforce that reserve requirements are not determined by MAPE alone: models with similar average accuracy can imply very different one-sided tail requirements (Reserve) under under-forecast risk. Weather integration narrows the error distribution across utilities (Fig. 3), which can translate directly into reduced upward reserve needs during temperature-driven demand spikes.
7.3 Computational Efficiency
The performance of Mamba-based models can be attributed to three factors: (1) selective state spaces capture long-range dependencies, (2) complexity enables extended 240h context windows, and (3) parameter efficiency. In our implementations, PowerMamba is particularly compact (2.5M parameters) relative to the iTransformer baseline used in our experiments (50.0M), enabling lower-latency inference and reduced memory footprint.
7.4 Foundation Model Performance
@@
7.5 Limitations from Unobserved BTM Generation
A recurring theme in our results is the ceiling imposed by data observability. Our "grey box" experiments with static NEMs registry features (Section 6.5) yielded only marginal gains. Figure 7 shows that even with capacity data, the model struggles to capture the dynamic, weather-dependent ramping of distributed solar. This points to a fundamental limitation: net load signals conflate grid demand with behind-the-meter (BTM) generation, and static capacity metadata is insufficient to disentangle them.
@@ 7.5 Limitations from Unobserved BTM Generation
Baseline Coverage. Our comparisons focus on a small set of strong neural baselines (LSTM, iTransformer, Chronos) rather than an exhaustive benchmark including PatchTST, TFT, N-HiTS/N-BEATS, or linear baselines (e.g., DLinear/NLinear). Future work should include autocorrelation-aware significance testing for a more complete statistical analysis.
8 Conclusion
Mamba-based state space models achieve competitive accuracy for California grid load forecasting. With weather integration, PowerMamba achieves 4.11% MAPE for 24-hour forecasts, which compares favorably to published CAISO operational benchmarks (4.55% MAPE; contextual reference). Critically, our walk-forward evaluation demonstrates that operational tail risk (Reserve) can diverge materially from average accuracy, motivating grid-specific reliability metrics alongside MAPE for safety-critical deployment decisions.
Weather Integration Across Regions. Weather integration improves 24-hour MAPE for several TAC areas (Appendix Table 7), particularly smaller and more volatile systems. Future work should evaluate tail-risk improvements under a strictly matched evaluation set (same timestamps across models).
Probabilistic calibration must be paired with bias controls. Multi-quantile objectives can reduce large-error events and tail reserve requirements, but can also be “gamed” by systematic inflation if Bias/OPR are not monitored and constrained. Our Bias/OPR-constrained objective provides an operator-legible way to enforce auditable trade-offs between tail-risk and schedule inflation (Table 2, Table 3). Operational Deployment Implications. Reduced tail under-forecast risk (Reserve) can translate to lower upward reserve requirements and reduced reliance on fast-ramping resources. The computational efficiency of Mamba architectures enables real-time deployment without significant infrastructure upgrades, while calibrated uncertainty quantification from Mamba-ProbTSF supports probabilistic reserve scheduling.

### v6

@@ IV-D Bias-Constrained Probabilistic Objective
Practically, we use these constrained objectives during fine-tuning to learn a calibrated distribution while preventing trivial “safety” via inflation. We then use (via ) to select an operationally conservative schedule or reserve policy from the output distribution. Appendix F describes how we empirically estimate from CAISO market data [35].
The Operator Evaluation Triad. To synthesize these competing operational objectives and ensure accountability, we propose evaluating reliability using a three-part diagnostic triad alongside standard MAPE:
1.
Reserve and UPR: Quantifies true tail risk (e.g., the upward generation capacity required to cover 99.5% of under-forecast events).
2.
Bias24h: Audits severity of systematic schedule inflation.
3.
OPR (Over-Prediction Rate): Audits frequency of schedule inflation.
Reporting this triad ensures that apparent improvements in tail risk (Reserve) are not simply “fake safety” achieved by shifting the mean bias upward. The constraints and explicitly enforce these trade-offs during training (evaluated in Section VI).
IV-E NEMs (BTM Registry) Feature Integration
While weather covariates proxy physical forcing, utility planning datasets often contain static NEMs/registry-derived BTM features (e.g., installed PV and storage capacity by region). We developed an architectural adaptation to fuse these static metadata on top of the weather-aligned baseline to test whether they provide incremental predictive value.
How NEMs features are integrated. NEMs/registry features are static or slow-moving, so we integrate them as an additional BTM input stream. To avoid physically irrelevant feature injection (e.g., PV capacity “explaining” nighttime net load), we apply a simple daylight prior (implemented as a time-of-day gate). When weather inputs are present, we further apply a lightweight context-conditioned modulation so the BTM contribution can vary smoothly by regime (e.g., daylight/irradiance conditions) rather than acting as an unconditional bias.
then fuse additively before the sequence encoder:
Figure 2 summarizes this “weather + NEMs” fusion path.
The incremental impact of this feature integration on walk-forward performance is evaluated in Section VI-E.
V Experimental Setup
V-A Dataset
@@
Preprocessing. Load values are normalized using z-score standardization computed only on training data to prevent information leakage. Temporal features include hour-of-day (0–23) and day-of-week (0–6), encoded as sinusoidal embeddings following standard practice [6]. For weather-integrated models, we include 8 meteorological covariates with thermal lag alignment, as described in Section V-C.
V-B CAISO Market Context and Benchmarks
@@
V-C Thermal Lag Parameter Determination
@@
Evaluation Protocol (Walk-Forward Backtest). We evaluate using rolling-origin walk-forward backtesting over Nov 2023–Nov 2025. Starting with an initial 180-day training history, we refit every 90 days using an expanding training window; each refit uses a 30-day validation window immediately preceding the cutoff for early stopping. We then evaluate one forecast per day (stride 24h) over the subsequent block until the next cutoff (48h horizon). All reported metrics in Tables I–II are computed on the concatenated set of walk-forward timestamps (477 evaluation windows for CA ISO-TAC).
V-D Models Under Comparison
We evaluate five neural architectures against an additional zero-shot foundation model baseline, enabling direct comparison across parameter efficiency and accuracy.
State Space Models.
@@ V-D Models Under Comparison
Transformer and Recurrent Baselines.
PatchTST [27]: Channel-independent Transformer (2.0M parameters), segmenting univariate channels into fixed-length patches. Represents the strongest load-only architecture in recent multi-ISO benchmarks.
iTransformer [28]: Inverted Transformer (6.5M parameters at ), tokenizing variables for cross-variate attention.
LSTM [37]: 2-layer bidirectional LSTM (2.6M parameters), representing the industry-standard recurrent approach.
Chronos [33]: Foundation models evaluated in the zero-shot regime. We primarily evaluate the 8M-parameter Small variant, and include the 200M-parameter Base variant in our discussion of scaling laws.
This comparison tests whether Mamba’s efficiency can match the accuracy of Transformer models, and whether the channel-independent (PatchTST) or cross-variate (iTransformer) paradigm better utilizes weather covariates for grid forecasting. PowerMamba (2.5M parameters) is particularly compact.
V-E Model Hyperparameters
Following the experimental protocol of PowerMamba [9], all Mamba variants share consistent architectural hyperparameters: (embedding dimension), (state dimension), (convolution kernel), expansion factor 2, and 2 bidirectional encoder layers. PatchTST uses with patch length ; iTransformer uses ; LSTM uses with 2 bidirectional layers. Details are provided in Appendix B. Weather integration adds 0.3M parameters for SSMs/LSTM and 0.7M for PatchTST.
V-F Training Protocol
@@
Fixed-split fine-tuning (loss-function ablation). For the loss-function ablation and bias/OPR constraints (Table II), we fine-tune weather-initialized checkpoints on a fixed split using the trainer framework: AdamW (lr ) with OneCycleLR, up to 150 epochs with patience 60, and multi-quantile heads (, weights ). Bias/OPR constraints are applied at the 24h lead time (; index 23 if using 0-based arrays for a 48h horizon forecast).
Reproducibility. All experiments run on NVIDIA RTX 5090 GPUs. We report the mean and standard deviation across three random seeds ({42, 123, 456}) for our main architecture experiments. Additional robustness details are in Appendix C.
V-G Evaluation Metrics
@@ V-H Statistical Analysis
VI Results
VI-A Multi-Horizon Performance (walk-forward)
Evaluation regime. This subsection reports results from the rolling-origin walk-forward backtest described in Section V-C.
Table I presents multi-horizon MAPE. PowerMamba is the strongest Mamba variant at the 1-hour horizon, while iTransformer achieves the lowest MAPE at longer horizons (6h–24h) in load-only evaluation. Crucially, as the rightmost column demonstrates, explicit weather integration substantially improves 24-hour accuracy across all comparable architectures, yielding the best overall performance (PowerMamba at 3.68%). These accuracy rankings, however, do not fully determine operational tail risk (Reserve), motivating the grid-specific metrics reported later.
| Model | Params | 1h | 6h | 12h | 24h (Base) | 24h (Weather) |
| S-Mamba | 16.4M | 5.49 0.12 | 3.86 0.02 | 3.40 0.47 | 8.03 0.18 | 4.47 |
| Mamba-ProbTSF | 16.4M | 5.23 0.04 | 3.91 0.19 | 3.30 0.08 | 7.89 0.17 | 4.52 |
| PowerMamba | 2.5M | 4.44 0.20 | 4.10 0.09 | 3.62 0.30 | 7.31 0.06 | 3.68 |
| PatchTST‡ | 2.0M | — | — | — | 4.89 | 4.72 |
| LSTM | 2.6M | 7.85 0.20 | 5.24 0.13 | 9.46 0.32 | 9.99 0.13 | —§ |
| iTransformer | 6.5M | 4.64 0.10 | 3.36 0.05 | 3.10 0.21 | 6.85 0.08 | 4.15 |
| Chronos (zero-shot)† | 8.0M | 2.10 | 3.21 | 3.72 | 7.59 | —§ |
| †Zero-shot evaluation; no training or uncertainty estimation. | ||||||
| ‡PatchTST was specifically evaluated at the 24h horizon to match multi-ISO benchmarks. | ||||||
| §LSTM and Chronos lack tailored weather integration mechanisms in our current implementation. | ||||||
VI-B Error Analysis: The Impact of Weather on Forecast Reliability (walk-forward)
Evaluation regime. This subsection uses the walk-forward backtest forecasts (all lead times within the 48h horizon) described in Section V-C.
@@ VI-B Error Analysis: The Impact of Weather on Forecast Reliability (walk-forward)
VI-C Weather Integration Results (walk-forward)
Evaluation regime. This subsection reports walk-forward results on matched evaluation windows (Section V-C).
Based on the error analysis, we integrated thermal-lagged weather covariates into the architectures. As shown in Figure 4, weather integration changes the forecast error distribution; improvements are not uniform across architectures and regions and should be reported on matched evaluation windows. Comprehensive 24-hour MAPE results across all TAC areas are provided in Appendix Table VI.
@@ VI-C Weather Integration Results (walk-forward)
VI-D Probabilistic Calibration and Bias/OPR-Constrained Objectives (fixed split)
Evaluation regime. This subsection reports fixed-split fine-tuning results for loss-function ablations (Section V-F); it is not directly comparable to the walk-forward results above.
@@ VI-D Probabilistic Calibration and Bias/OPR-Constrained Objectives (fixed split)
Table II presents an operator-legible breakdown (CA ISO-TAC, 24h), explicitly reporting Bias24h and OPR alongside tail-risk Reserve. Three high-signal observations:
@@
Effect of probabilistic calibration on tail events. Figure 6 compares (i) Weather MSE, (ii) WeatherMulti-Q, and (iii) WeatherMulti-Q+BiasCtrl on a single consistent evaluation split (same targets across variants). Multi-quantile calibration reduces large-error counts relative to Weather MSE, while adding Bias/OPR constraints partially trades off tail improvements for reduced inflation risk (Table II).
VI-E NEMs (BTM Registry) Integration Results
Evaluation regime. This subsection reports walk-forward results on utilities where NEMs/registry features are available (PGE, SCE, SDGE, TIDC) using the fusion architecture described in Section IV-E.
@@ VI-E NEMs (BTM Registry) Integration Results
VII Discussion
VII-A Performance Comparison with CAISO Benchmarks
Table III provides contextual comparison with published CAISO operational forecasts and commercial alternatives.
@@ VII-A Performance Comparison with CAISO Benchmarks
| LSTM Baseline | 6.49% | +42.7% | 2.6M | No |
| PatchTST | 4.89% | +7.5% | 2.0M | No |
| iTransformer | 5.37% | +18.0% | 6.5M | No |
@@
| iTransformer + Weather | 4.15% | -8.8% | 6.5M | Yes |
| PatchTST + Weather | 4.72% | +3.7% | 2.0M | Yes |
@@
The performance gains are particularly significant given the model efficiency: PowerMamba achieves competitive accuracy relative to published CAISO benchmarks with only 2.5M parameters. Among the Transformer baselines, PatchTST achieves the strongest load-only accuracy (4.89%) with only 2.0M parameters, confirming its suitability for load-only deployment.
VII-B Channel-Independent vs. Cross-Variate Attention
The load-only vs. weather-augmented ranking reversal traces to a fundamental architectural difference. iTransformer’s symmetric variate tokenization treats weather features as first-class tokens in the same attention space as load, enabling fine-grained load–weather interaction at each time step. PatchTST processes channels independently, so weather information can only influence load prediction indirectly through cross-attention sublayers. This division explains why iTransformer achieves a larger weather benefit than PatchTST on CAISO, and why the gap widens on weather-sensitive grids in our companion multi-ISO benchmark. For practitioners, this suggests deploying PatchTST for load-only forecasting, and switching to iTransformer or weather-aware SSM variants when weather forecasts are available.
VII-C Temperature, Tail Risk, and Operator Implications
Our error analysis shows a statistically significant but modest association between temperature and forecast error magnitude (), supporting the operational intuition that weather is a relevant covariate and that loss-function modifications cannot substitute for missing exogenous information. More importantly for operations, the walk-forward results reinforce that reserve requirements are not determined by MAPE alone: models with similar average accuracy can imply substantially different one-sided tail requirements (Reserve) under under-forecast risk. Weather integration narrows the error distribution across utilities (Fig. 4), which can translate directly into reduced upward reserve needs during temperature-driven demand spikes.
VII-D Computational Efficiency
The performance of Mamba-based models can be attributed to three factors: (1) selective state spaces capture long-range dependencies, (2) complexity enables extended 240h context windows, and (3) parameter efficiency. In our implementations, PowerMamba is particularly compact (2.5M parameters) relative to the iTransformer baseline used in our experiments (6.5M), enabling lower-latency inference and reduced memory footprint.
VII-E Foundation Model Performance
@@
VII-F Limitations from Unobserved BTM Generation
A recurring theme in our results is the ceiling imposed by data observability. Our "grey box" experiments with static NEMs registry features (Section IV-E) yielded only marginal gains. Figure 7 shows that even with capacity data, the model struggles to capture the dynamic, weather-dependent ramping of distributed solar. This points to a fundamental limitation: net load signals conflate grid demand with behind-the-meter (BTM) generation, and static capacity metadata is insufficient to disentangle them.
@@ VII-F Limitations from Unobserved BTM Generation
Baseline Coverage. Our CAISO-focused evaluation is complemented by a companion multi-ISO benchmark [38] covering six US grid operators, five forecast horizons, and multitype generalization to solar, wind, and price forecasting. Future work should include autocorrelation-aware significance testing for a more complete statistical analysis.
VIII Conclusion
Mamba-based state space models achieve competitive accuracy for California grid load forecasting. With weather integration, PowerMamba achieves 3.68% MAPE for 24-hour forecasts, which compares favorably to published CAISO operational benchmarks (4.55% MAPE; contextual reference). Critically, our walk-forward evaluation demonstrates that operational tail risk (Reserve) can diverge materially from average accuracy, motivating grid-specific reliability metrics alongside MAPE for safety-critical deployment decisions.
Weather Integration Across Regions. Weather integration improves 24-hour MAPE for several TAC areas (Appendix Table VI), particularly smaller and more volatile systems. Future work should evaluate tail-risk improvements under a strictly matched evaluation set (same timestamps across models).
Probabilistic calibration must be paired with bias controls. Multi-quantile objectives can reduce large-error events and tail reserve requirements, but can also be “gamed” by systematic inflation if Bias/OPR are not monitored and constrained. Our Bias/OPR-constrained objective provides an operator-legible way to enforce auditable trade-offs between tail-risk and schedule inflation (Table II). Operational Deployment Implications. Reduced tail under-forecast risk (Reserve) can translate to lower upward reserve requirements and reduced reliance on fast-ramping resources. The computational efficiency of Mamba architectures enables real-time deployment without significant infrastructure upgrades, while calibrated uncertainty quantification from Mamba-ProbTSF supports probabilistic reserve scheduling.